OpenStudy (anonymous):

use the method of elimination to solve the system of linear equations a+8b=10 a+4b=2

OpenStudy (anonymous):

a+8b-a-4b=10-2, then 4b=8 so b=2, a+8=2 then a =-6

OpenStudy (abb0t):

wut?!

OpenStudy (anonymous):

make both equal to zero a+8b-10=0.....(1) a+4b-2=0....(2) subtract (1) (2) 4b-8=0 4b=8 b=2 sub. in original equation (1) a+8b=10 a+8(2)=10 a=-16+10 a=-6

OpenStudy (anonymous):

thnks @kyusakazi and @physicsboffin

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

@Avon u're welcome